A Delta 727 travelled 560mph with the wind and 480mph against the wind. Find the speed of the plane in still air and the speed of the wind.
Answer by mananth(16949)

let speed of plane in still air be x mph
speed of wind be y mph
time = d/r
with wind time
560/(x+y)=1
x+y =560........................1
...
against wind time
480/(x-y)=1
x-y =480........................2
add the two equations (1)&(2)
2x=1040
/2
x=1040/2
x= 520 mph speed of plane in still air
y= 40 mph speed of wind
...
m.ananth@hotmail.ca